OpenStudy (anonymous):

Please help!! Its important! Solve using substitution! A jet boat is used to take canoers 200 miles upriver to their starting point. As the driver of the jet boat, Jeremy knows that it will take him 5 hours to get the canoers upstream, but only 4 hours to get the boat back downstream after dropping the canoers off, as long as he keeps the boat moving at a constant speed. How fast would the boat be moving in still water? A. 5 mph B. 40 mph C. 45 mph D. 50 mph

OpenStudy (neer2890):

Let speed of boat in still water is x miles/hr and speed of river is y miles/hr. so upstream, speed of boat=(x-y) miles/hr downstream, speed of boat=(x+y) miles/hr

OpenStudy (neer2890):

now. we know Distance= speed*time so upstream, 200=(x-y)5 i.e., 5x-5y=200 Downstream, 200=(x+y)4 i.e., 4x+4y=200 now solve the two equations by any method you know to solve equations of two variables to find x and y. x is the speed of boat in still water. and y is the speed of stream or river.

OpenStudy (anonymous):

Thank you so much! This is so helpful. I was so stuck. Work problems aren't my thing.

OpenStudy (neer2890):

nvm... you're welcome..:)

OpenStudy (anonymous):

No no, continue. I still need help. XD

OpenStudy (neer2890):

5x-5y=200......(i) 4x+4y=200......(ii) multiply (i) with 4 and (ii) by 5 and add both.

OpenStudy (neer2890):



OpenStudy (anonymous):

u can also take the 200 miles divide it by 5 hrs and get 40 (the going trip) which is mph then divide 200 miles by 4 hrs and get 50 (the return trip)mph the add 40+50 =90mph for the total trip and divide by 2 to get the average of 45mph

OpenStudy (anonymous):

the way @neer2890 did it is correct also!

OpenStudy (neer2890):

yeah... but adopt the method used by @sneakergyrl when you understand the previous one.. because it's a shortcut.